# The dynamics of the Reddit collective action leading to the GameStop short squeeze

### v3

@@ Conclusion
In January 2021, the activity of WSB users anticipated financial trends. The high temporal resolution of our analysis enabled us to identify three distinct phases of online behavior and engagement with the stock market, characterized by Discussion, Action and Visibility, respectively. Our results support the hypothesis that the WSB community coordinated a collective financial strategy, which then had a large follow up on other social networks as well. First, we observed that the strong anticipatory relation between WSB users activity and changes in GME trading volume emerged sharply on January 13th. This suggests a sudden transition between a phase characterized by Discussion and one characterized by financial Action. Second, we found that – at around the same time – the value of the collective position on GME started to closely mirror the evolution of the market capitalization of GME. We estimated that the size of the total position of WSB users amounted to at least 1% of GME market capitalization, demonstrating a tangible engagement of the WSB community in the stock market. This suggests that other market players joined the short squeeze after it was initiated by WSB users, despite having access to information about GME that was presumably not inferior to that of the WSB users[55]. After the price hike of GME and a Tweet by Elon Musk on January 27th, the short squeeze became widely known, and a broader public started to follow Reddit conversations.
@@
Our study comes with limitations. First, our findings refer to a specific case study, the GME short squeeze, which so far represents a unique event due to its wide resonance. If and when such events will occur again, we could understand whether the patterns we have observed can be generalized to different contexts. Secondly, our analysis primarily relies on Granger test, a standard statistical framework that yet relies on linear models – which may oversimplify complex relationships. Recognizing this limitation, we have explored nonlinear techniques; notably, an analysis based on Convergent Cross-Mapping [52] corroborates our main results (see Supplementary Materials, Section 5). Third, due to limitation of the available data, our analysis focused solely on Reddit and Twitter, neglecting offline interactions or happening on other online platforms (e.g., Telegram or Discord), as well as the actions of institutional investors. Finally, the proxy used to estimate the size of WSB collective position on GME, specifically from user screenshots, are inherently partial and noisy.
@@
Given the limitations of our approach, we provide extensive evidence regarding the possible influence exerted by Reddit users on the GameStop short squeeze. However, determining a genuine causal link – that Reddit users directly fueled the price rally – is a ill-posed problem given the data we can access. Factors such as the true portfolio composition of each Reddit user and the potential psychological contagion within Reddit discussions add uncertainty for the attribution of direct causation.
@@
Overall, our results contribute to a deeper understanding of how digital platforms can catalyze collective action and influence real-world outcomes. In an era characterized by rapid technological advancements, the pace of our daily lives is accelerated to an unprecedented rate [56]. This phenomenon has also impacted the financial landscape, where the emergence of commission-free trading platforms started a new age of dynamism and accessibility [31]. Our findings highlight the importance of investigating the dynamic interplay between online engagement and financial market indicators at an hourly scale (or even finer) [13, 14, 10, 12]. Hence, they suggest that strategies for monitoring and, eventually, halting social media-driven movements need to be adaptable and responsive to these evolving dynamics.
Methods
@@ 0.1 Data Descriptions
We gathered Reddit conversation data from Pushshift [57], an API that regularly copies activity data of Reddit. We queried the service to retrieve information about posts and comments on WSB from December 01, 2020 to July 1, 2021. The dataset was cleaned by removing posts/comments by Reddit bots (see Supplementary Materials, Section 1) as well as by "[deleted]" users, who deleted their account at the time Pushshift API collected data. To obtain the occurrences of a given stock within WSB conversations, we counted how many times the ticker symbol of the stock (e.g., ‘NOK’ for Nokia) appears as a regular expression in the raw text of a post/comment [41].
Within the WSB community, a common practice consists in posting financial investments by sharing screenshots of open positions, primarily gains, losses, or orders. These “committed” screenshots were identified and classified using computer vision techniques by Lucchini et al.[40]. Using Tesseract, an open-source optical character recognition [58], we extract textual information from these screenshots, converting image-based text into editable and searchable content. We focused on screenshots containing keywords such as ‘gme,’ ‘gamestop,’ ‘investing,’ and ‘value,’ identified through manual inspection, and subsequently narrowed down the sample to 6525 screenshots from November 1, 2020 to February 04, 2021. After parsing the structures, we isolated dollar-sign-prefixed numbers, determining the final value of the position as the highest extracted value. However, such value may represent trading volume or the price of GME taken from online trading applications, rather than an actually investment. To address this, we filtered out screenshots with values below $10^{2}$ and above $10^{6}$, as these are likely indicative of price or trading volume. We manually examined values above $10^{6}$ to validate our findings. Additionally, we manually validated 1000 screenshots and determined that this procedure extracts the true value with an accuracy of 0.85 and a root mean squared error (RMSE) of 0.47. The RMSE is computed using log-transformed data, and accuracy is based on a binary classification problem, where correctness is determined by applying a threshold of 0.05 to the absolute logarithmic ratio between the extracted and true values (see Supplementary Materials, Section 2).
@@ 0.3 Granger test and Multivariate Vector Autoregressive Model
where the first equation is the restricted model and the second the full model. After fitting these models, the p-values are computed from the F-statistic, which reflects the difference in the residual sum of squares ($RSS$). These p-values indicate the likelihood of obtaining results as extreme as the observed F-statistic under the assumption that the restricted model is correct. The Granger Index, on the other hand, focuses only on the predictive accuracy of the full model in relation to the restricted model, defined as $GI=\ln(RSS_{restricted}/RSS_{full})$. To conduct the Granger test, we need first to evaluate the stationarity of signals within the specified windows. The results of the Augmented Dickey-Fuller test are reported in Supplementary Materials, Section 4.
Following the full model of the Granger test, we consider a Multivariate Vector Autoregressive model to analyze relationships and dependencies among two or more time series simultaneously. For instance, within a specified model aimed at forecasting variable $y$ by using historical data at a certain lag, denoted as $y(t-l)$ and $x(t-l)$, we derive the coefficients $c^{(y)}_{l}$ and $c^{(x)}_{l}$. These coefficients capture the respective influences and contributions of the lagged values of $y$ and $x$ in predicting the future values of the variable $y$. The significance of a particular coefficient is established by evaluating its corresponding p-value, computed similarly to the Granger analysis. Both for Granger and Multivariate Vector Autoregressive Model, we analyzed the data by sliding it within 15-day windows, shifting by one day each time. The resulting length of the time series is thus of 120 points (8 hourly points per day) – a reasonable number to perform linear regression analysis [60, 61, 62].
Acknowledgments
A.D. acknowledges the helpful discussions on Twitter and screenshot data analysis with Lorenzo Lucchini. A.D. acknowledges the helpful discussions on causality and Reddit conversations analysis with Riccardo Di Clemente, Anna Mancini and Gabriele Di Antonio. A.D. acknowledges Danish Data Science Academy for their funding support ("Grant ID DDSA-V-2023-008"). L.M.A. acknowledges support from the Carlsberg Foundation through the COCOONS project (CF21-0432). G.C. acknowledges support from “Deep ’N Rec” Progetto di Ricerca di Ateneo of University of Rome Tor Vergata.
@@ Acknowledgments
Supplementary Materials, Section 2 : Screenshot analysis
As detailed in the main text, we employ computer vision techniques to extract actual financial investments within the WSB community. In order to provide proof of one’s financial position, users in the WSB community typically share screenshots of their open positions. Lucchini et al.[40] quantified the WSB community’s commitment to GME by parsing these screenshots using a combination of text analysis and computer vision techniques. To extract the collective investments we rely on this previous analysis and consider the dataset of 78468 screenshots as our initial dataset. We note that these screenshots exhibit arbitrary structures, as WSB users do not use a specific app and screenshots are handmade, resulting in a lack of standardized format. Through manual inspection of a sample of the screenshots, we identify a set of keywords that prove useful in narrowing down the selection to those representing actual investments (6525 screenshots), as discussed in the main text. Our extraction technique, for instance, does not accommodate the possibility of multiple orders in a single screenshot, as we extract only the highest value.
Supplementary Figure S1: Comparison between OCR and human extraction and stocks rank. a) Scatter plot displaying the comparison between investments extracted using OCR and human visual inspection. b) Treemap rank chart showing the occurrences of stock tickers in screenshots.
Stock A Stock B Number of Screenshots
Stock A Stock B Number of Screenshots
AMC PLTR 56
GME NIO 62
NOK PLTR 62
NIO PLTR 65
AMC NIO 70
NIO TSLA 80
PLTR TSLA 81
AMC TSLA 98
BB PLTR 132
GME TSLA 148
GME PLTR 183
BB NOK 457
GME NOK 484
AMC BB 648
AMC NOK 698
BB GME 719
AMC GME 1773
@@
As discussed in the main text, we perform the detrended cross-correlation analysis on BB, AMC, NOK. However, in this case, our analysis focuses on data spanning from January 31st to July 1st, 2021. This time frame is chosen as, prior to January, these stocks were not extensively discussed, as reported in the main text. We observe that, unlike the GME case reported in the main text, the cross-correlation coefficients exhibit variation and dependence on the scale $t$ (see Figure S2). This variability could be attributed to the limited number of considered time scales and it is similar to the variability observed by Podobnik et al.[59] for the New York and Shanghai Stock Exchange Composite index. Nevertheless, the power-law relationship of the covariances and variances (see Figure S3), as well as the fitted scale exponent reported in Table 7, supports the findings reported in the main text.
@@
Supplementary Figure S2: Detrended Cross Correlation coefficients for the following combinations of signals: price / trading volume (light blue), price / Reddit activity (pink), trading volume / Reddit activity (purple). In panel a, BB, in panel b, AMC, and in panel c, NOK.
Supplementary Figure S3: Detrended Variances for each signal: price (red), trading volume (blue) and Reddit activity (green); detrended covariances for their paired combinations (same color-scheme of Figure S2). In panel a, BB, in panel b, AMC, and in panel c, NOK.
We perform an additional analysis by examining the detrended variance of each signal removing the mean of the signal in each window to further characterise the non-linear dependencies. This latter exercise allows us to estimate the Hurst exponent [51, 63]. In this case, instead of considering the variance as a function of the window $t$, we examine the rescaled range $R/S$ [51]. The rescaled range is the ratio between the range of the signal, which is the difference between the maximum and minimum values in the cumulative sum of the residuals from the mean, and the standard deviation of the residuals (variance up to a square root). We consistently observe similar results across the analyzed stocks: the Reddit signals exhibit an almost power-law behavior in the rescaled range as a function of the window $t$, along with a Hurst exponent greater than 0.5 (see Figure S4 and Table 8). In this case, we represent all windows $t$ since the fluctuations, though less pronounced than the previous analysis, align with common findings in the literature [51]. These results suggest a persistent and trending behavior in the Reddit changes, indicating the presence of long-range dependencies in the signal. To accurately determine the Hurst exponent or the scale exponent, it is recommended to include a broader range of time scales, such as years, in the analysis [51]. This analysis serves as an exercise in exploring possible non-linear dependencies in social media / financial markets dynamics.
@@
Supplementary Table 8: Hurst exponent for all stocks and signals. The exponents are fitted starting from window 10 to 200.
@@
Date t (Twitter) p-value (Twitter) t (Trading Volume) p-value (Trading Volume) t (Reddit) p-value (Reddit) 2020-12-17 -4.296 $0$4.299 $0$4.429 0.000 2020-12-18 -4.251 0.001 -4.236 0.001 -4.504 0.000 2020-12-19 -4.894 $0$4.233 0.001 -4.405 0.000 2020-12-20 -4.894 $0$4.233 0.001 -4.405 0.000 2020-12-21 -4.894 $0$4.233 0.001 -4.405 0.000 $-2$5.000 $0$4.168 0.001 -4.350 0.000 2020-12-23 -4.627 $0$5.939 $0$5.031 0.000 2020-12-24 -5.182 $0$4.987 $0$5.315 0.000 2020-12-25 -5.149 $0$4.134 0.001 -4.947 0.000 2020-12-26 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-27 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-28 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-29 -4.751 $0$3.880 0.002 -4.666 0.000 2020-12-30 -4.729 $0$4.091 0.001 -4.712 0.000 2020-12-31 -1.841 0.360 -4.280 $0$5.348 0.000 2021-01-01 -3.202 0.020 -4.657 $0$5.688 0.000 2021-01-02 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-03 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-04 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-05 -2.938 0.041 -4.258 0.001 -5.288 0.000 2021-01-06 -2.698 0.074 -4.552 $0$4.940 0.000 2021-01-07 -2.715 0.072 -4.689 $0$4.820 0.000 2021-01-08 -2.788 0.060 -5.738 $0$5.764 0.000 2021-01-09 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-10 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-11 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-12 -3.698 0.004 -6.108 $0$5.725 0.000 2021-01-13 -3.815 0.003 -5.685 $0$4.988 0.000 2021-01-14 -1.185 0.680 -4.290 $0$4.738 0.000 2021-01-15 -5.310 $0$4.294 $0$4.785 0.000 2021-01-16 -5.680 $0$4.425 $0$5.053 0.000 2021-01-17 -5.680 $0$4.425 $0$5.053 0.000 2021-01-18 -5.680 $0$4.425 $0$5.053 0.000 2021-01-19 -5.346 $0$4.167 0.001 -4.937 0.000 2021-01-20 -5.225 $0$4.147 0.001 -4.752 0.000 2021-01-21 -5.104 $0$3.747 0.004 -4.792 0.000 $-2$5.362 $0$4.006 0.001 -4.843 0.000 2021-01-23 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-24 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-25 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-26 -4.094 0.001 -4.004 0.001 -5.068 0.000 2021-01-27 -4.972 $0$3.601 0.006 -10.878 0.000 2021-01-28 -4.298 $0$4.008 0.001 -3.058 0.030 2021-01-29 -4.189 0.001 -4.101 0.001 -2.560 0.102 2021-01-30 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-01-31 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-02-01 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-02-02 -3.254 0.017 -4.352 $0$2.881 0.048 2021-02-03 -3.864 0.002 -4.358 $0$2.767 0.063 2021-02-04 -3.460 0.009 -3.839 0.003 -2.696 0.075 2021-02-05 -3.702 0.004 -3.995 0.001 -2.989 0.036 2021-02-06 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-07 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-08 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-09 -3.001 0.035 -3.754 0.003 -2.262 0.184 2021-02-10 -2.733 0.069 -3.412 0.011 -2.268 0.182 2021-02-11 -2.702 0.074 -3.446 0.009 -2.765 0.063 2021-02-12 -3.432 0.010 -3.196 0.020 -2.123 0.235 2021-02-13 -3.045 0.031 -3.006 0.034 -1.854 0.354

### v4

@@ 3 Discussion
In January 2021, the activity of WSB users anticipated financial trends. The high temporal resolution of our analysis enabled us to identify three distinct phases of online behavior and engagement with the stock market, characterized by Discussion, Action and Visibility, respectively. First, we observed that the strong anticipatory relation between WSB users activity and changes in GME trading volume emerged sharply on January 13th. This suggests a sudden transition between a phase characterized by Discussion and one characterized by financial Action. Second, we found that – at around the same time – the value of the collective position on GME started to closely mirror the evolution of the market capitalization of GME. We estimated that the size of the total position of WSB users amounted to at least 1% of GME market capitalization, demonstrating a tangible engagement of the WSB community in the stock market. This suggests that other market players joined the short squeeze after it was initiated by WSB users, despite having access to information about GME that was presumably not inferior to that of the WSB users [45]. After the price hike of GME and a Tweet by Elon Musk on January 27th, the short squeeze became widely known, and a broader public started to follow Reddit conversations.
@@
Our study comes with limitations. First, our findings refer to a specific case study, the GME short squeeze, which so far represents a unique event due to its wide resonance. If and when such events will occur again, we could understand whether the patterns we have observed can be generalized to different contexts. Secondly, our analysis primarily relies on Granger test, a standard statistical framework that relies on linear models. As such, it may oversimplify complex relationships, such as those involved in collective action between social networks and financial markets, where feedback loops and non-linear dynamics can occur. Recognizing this limitation, we have explored nonlinear techniques; notably, an analysis based on Convergent Cross-Mapping [41] corroborates our main results (see Supplementary Materials, Section 3). Third, due to limitation of the available data, our analysis focused solely on Reddit and Twitter, neglecting offline interactions or happening on other online platforms (e.g., Telegram or Discord), as well as the online/offline actions of institutional investors. Finally, the proxy used to estimate the size of WSB collective position on GME, specifically from user screenshots, are inherently partial and noisy. We also have no information on the identities of the investors behind the screenshots, so we cannot exclude the potential interference by institutional investors in the financial discourse on Reddit.
@@
Despite such limitations, we have provided extensive evidence regarding the possible influence exerted by Reddit users on the GameStop short squeeze. However, determining a genuine causal link – that Reddit users directly fueled the price rally – is not possible with the data we can access. Factors such as the true portfolio composition of each Reddit user and the potential psychological contagion within Reddit discussions add uncertainty for the attribution of direct causation.
@@
Overall, our results add further evidence that digital platforms can catalyze collective action and influence real-world outcomes. In an era characterized by rapid technological advancements, the pace of our daily lives is accelerated to an unprecedented rate [46]. This phenomenon has also impacted the financial landscape, where the emergence of commission-free trading platforms started a new age of dynamism and accessibility [29]. Our findings highlight the importance of investigating the dynamic interplay between online engagement and stock market movements at an hourly scale (or even finer) [10, 11, 7, 9], as the coordinated action of retail traders through social media has become a no-longer-negligible market player.
4 Methods
@@ 4.1 Data Descriptions
We gathered Reddit conversation data from Pushshift [47], an API that regularly copies activity data of Reddit. We queried the service to retrieve information about posts and comments on WSB from December 01, 2020 to July 1, 2021. The dataset was cleaned by removing posts/comments by Reddit bots (see Supplementary Materials, Section 1) as well as by ”[deleted]” users, who deleted their account at the time Pushshift API collected data. To obtain the occurrences of a given stock within WSB conversations, we counted how many times the ticker symbol of the stock (e.g., ‘NOK’ for Nokia) appears as a regular expression in the raw text of a post/comment [26].
Within the WSB community, a common practice consists in posting financial investments by sharing screenshots of open positions, primarily gains, losses, or orders. These “committed” screenshots were identified and classified using computer vision techniques by Lucchini et al.[30]. Using Tesseract, an open-source optical character recognition [48], we extract textual information from these screenshots, converting image-based text into editable and searchable content. We focused on screenshots containing keywords such as ‘gme,’ ‘gamestop,’ ‘investing,’ and ‘value,’ identified through manual inspection, and subsequently narrowed down the sample to 6525 screenshots from November 1, 2020 to February 04, 2021. After parsing the structures, we isolated dollar-sign-prefixed numbers, determining the final value of the position as the highest extracted value. However, such value may represent trading volume or the price of GME taken from online trading applications, rather than an actually investment. To address this, we filtered out screenshots with values below $10^{2}$ and above $10^{6}$, as these are likely indicative of price or trading volume. We manually examined values above $10^{6}$ to validate our findings. Additionally, we manually validated 1000 screenshots using log-transformed data, computing the logarithms of the ratios between true and estimated values. We determined that our procedure extracts the correct value with an accuracy of 0.85, where accuracy is based on a binary classifier where correctness is determined by applying a threshold of 0.05 to the absolute logarithmic ratios (see Supplementary Materials, Section 4). The root mean square error (RMSE) of the logarithmic ratios between true and estimated values is 0.47, and the point estimate (i.e., the average ratio for a single screenshot) amounts to -0.07. This indicates that the automated predictions slightly overestimate the actual values (on average +17%). However, this difference is more pronounced at screenshot values below $10^{3}$, with a relative logarithmic error of -0.1, while it approaches zero at higher values. Overall the automated extraction procedure leads to a total discrepancy of approximately +0.2% in the WSB posted collective position. Note that this validation was intended solely to assess the error of the employed framework in estimating the collective financial position – the incorrect estimates were not replaced. Indeed, missing posts, and possibly fake ones, represent a much larger source of error, which we cannot control.
@@ 4.3 Granger test and Multivariate Vector Autoregressive Model
where the first equation is the restricted model and the second the full model. After fitting these models, the p-values are computed from the F-statistic, which reflects the difference in the residual sum of squares ($RSS$). These p-values indicate the likelihood of obtaining results as extreme as the observed F-statistic under the assumption that the restricted model is correct. Size effects are given by the values of the regression coefficients $\widehat{a_{l}}$ and $\widehat{b_{l}}$, which capture the contribution of the lagged variables $y$ and $x$ in predicting $y$. The Granger Index, on the other hand, focuses only on the predictive accuracy of the full model in relation to the restricted model, defined as $GI=\ln(RSS_{restricted}/RSS_{full})$. To conduct the Granger test, we need first to evaluate the stationarity of signals within the specified windows. The results of the Augmented Dickey-Fuller test are reported in Supplementary Materials, Section 5.
Following the full model of the Granger test, we consider a Multivariate Vector Autoregressive model to analyze relationships and dependencies among two or more time series simultaneously. The significance of a particular regression coefficient is established by evaluating its corresponding p-value, computed similarly to the Granger analysis. Both for Granger and Multivariate Vector Autoregressive Model, we analyzed the data by sliding it within 15-day windows, shifting by one day each time. The resulting length of the time series is thus of 120 points (8 hourly points per day) – a reasonable statistics to perform linear regression analysis [50, 51, 43].
Declarations
  • Data Availability: Reddit conversation data used in this study (the same used by Mancini et al.[26] and Lucchini et al.[30]) can be retrieved from the Pushshift API at https://www.reddit.com/r/pushshift/ and were collected in December 2022. Twitter data used in this analysis are the same used by Lucchini et al.[30], and were retrieved from Twitter API before April 2022. Stock market data used in this analysis can be retrieved from the Polygon.io API at https://polygon.io and were collected in December 2022, while the GME Market Capitalization was collected in September 2023.
  • Code Availability: The data and code to reproduce the analysis is released on GitHub. For inquiries, please contact A.D. antde@dtu.dk.
  • Acknowledgements A.D. acknowledges the helpful discussions on Twitter and screenshot data analysis with Lorenzo Lucchini. A.D. acknowledges the helpful discussions on Reddit conversations analysis with Riccardo Di Clemente, Anna Mancini and Gabriele Di Antonio. A.D. acknowledges Danish Data Science Academy for their funding support (”Grant ID DDSA-V-2023-008”). L.M.A. acknowledges support from the Carlsberg Foundation through the COCOONS project (CF21-0432). G.C. acknowledges support from “Deep ’N Rec” Progetto di Ricerca di Ateneo of University of Rome Tor Vergata.
  • Author Contributions: A.D. and L.M.A gathered the data. A.D. performed the analysis and made the figures. L.A. and G.C designed the analysis and supervised the project. All authors discussed the results and contributed to the final manuscript.
  • Competing Interests: The authors declare no competing interests
@@
As discussed in the main text, we perform the detrended cross-correlation analysis on BB, AMC, NOK. However, in this case, our analysis focuses on data spanning from January 31st to July 1st, 2021. This time frame is chosen as these stocks were not extensively discussed prior to January, as reported in the main text. Unlike the GME case reported in the main text, we observe that the cross-correlation coefficients exhibit variation and dependence on the scale $t$ (see Figure S1). This variability could be attributed to the limited number of time scales considered and it is similar to the variability observed by Podobnik et al. [49] for the New York and Shanghai Stock Exchange Composite index. Nevertheless, the power-law relationship of the covariances and variances (see Figure S2), as well as the fitted scale exponent reported in Table 7, supports the findings reported in the main text.
@@
Supplementary Figure S1: Detrended Cross Correlation coefficients for the following combinations of signals: price / trading volume (light blue), price / Reddit activity (pink), trading volume / Reddit activity (purple), for BB (panel a), AMC (panel b) and NOK (panel c).
Supplementary Figure S2: Detrended Variances for each signal: price (red), trading volume (blue) and Reddit activity (green); detrended covariances for their paired combinations (same color-scheme of Figure S1); for BB (panel a), AMC (panel b) and NOK (panel c).
We perform an additional analysis by examining the detrended variance of each signal removing the mean of the signal in each window to further characterise the non-linear dependencies. This latter exercise allows us to estimate the Hurst exponent [40, 52]. In this case, instead of considering the variance as a function of the window $t$, we examine the rescaled range $R/S$ [40]. The rescaled range is the ratio between the range of the signal, which is the difference between the maximum and minimum values in the cumulative sum of the residuals from the mean, and the standard deviation of the residuals (variance up to a square root). We consistently observe similar results across the analyzed stocks: the Reddit signals exhibit an almost power-law behavior in the rescaled range as a function of the window $t$, along with a Hurst exponent greater than 0.5 (see Figure S3 and Table 8). In this case, we represent all windows $t$ since the fluctuations, though less pronounced than the previous analysis, align with common findings in the literature [40]. These results suggest a persistent and trending behavior in the Reddit changes, indicating the presence of long-range dependencies in the signal. To accurately determine the Hurst exponent or the scale exponent, it is recommended to include a broader range of time scales in the analysis – such as years [40]. Our analysis serves as an exercise in exploring possible non-linear dependencies in social media / financial markets dynamics.
@@
Supplementary Table 8: Hurst exponent for all stocks and signals. The exponents are fitted from window length 10 to 200.
@@
Supplementary Materials, Section 3: Convergent Cross-Mapping
To capture and quantify non-linear coupling relations between variables of the same dynamical systems, we employ a statistical technique known as Convergent Cross Mapping (CCM) [41]. Similar to Granger Test, the key concept behind CCM is to test whether the knowledge of the past states of variable $x$ can improve predictions of variable $y$ ’s future states. Unlike Granger Test, CCM does not assume that we can separate the information about $x$ from the rest of the systems, as it is often the case in nonlinear models [53, 54]. Given a time series $x(t)$, we consider the delay vectors $X_{\tau,D}(t)=(x(t),\dots,x(t-(D-1)\tau))$ and for large enough embedding dimension D (larger than the attractor dimensions) the manifold $M_{X}$ identified by $X$ reconstructs the original true manifold $M_{S}$ (or all the topological invariant of the dynamical systems $S$). Given two time series $x(t)$ and $y(t)$, we assume that these are generated from an unknown high-dimensional dynamical systems $S$. Using Takens’ theorem [55, 56, 57], we can construct from each time series $s$ a manifold $M_{s}$ and confront at a given time $t$ the points on one reconstructed manifold $(M_{X})$ with the other $(M_{Y})$. If $X\to Y$, then information from $X$ is stored in $Y$ and gets embedded in $M_{Y}$, thus it follows that we can use $M_{Y}$ to better predict $X$ (the manifold $M_{Y}$ contain the information of both $X$ and $Y$).
$\tilde{Y}_{\tau,D}(t)=(y(t),\dots,y(t-(D-1)\tau),x(t))$ and we perform a neighbor search using the first dimensions to predict last dimension, i.e., $x(t)$. The correlation $\rho_{YX}$ summarises the results of this prediction [41]. CCM requires two criteria to conclude synergistic relation: cross mapping and convergence [41]. Cross mapping is linked to the correlation $\rho$ and the stronger the coupling in one (or both) direction, the higher the value of $\rho$, since the prediction error using one manifold is either lower or equal to that of the other. Convergence refers to improving the cross-mapping with time-series length $T$ (sample size to reconstruct the manifold, referred as library size [41]) and it is related to Takens’ theorem. The more the data, the more precise is the reconstruction of the attractor, resulting in closer nearest neighbors and higher correlation $\rho$.
Supplementary Figure S5: Convergent Cross-Mapping for all the stocks. Correlation $(\rho)$ as a function of the length of time series $T$ for optimal lag $\tau$ and optimal dimension $D$ for different combinations of signals: Reddit-to-trading volume (orange, left panel), trading volume-to-Reddit (blue, left panel), trading volume-to-price (purple, center), price-to-trading volume (light green, center), price-to-Reddit (light blue, right panel), Reddit-to-price (pink, right panel). In panel a, BB $(\tau=6,D=5)$, in panel b, AMC $(\tau=6,D=7)$, in panel c, NOK $(\tau=6,D=6)$, and in panel d, GME $(\tau=7,D=5)$.
Supplementary Materials, Section 4: Screenshot analysis
As detailed in the main text, we employ computer vision techniques to extract actual financial investments within the WSB community. Indeed users in the WSB community typically share screenshots of their open positions. Lucchini et al. [30] quantified the WSB community’s commitment to GME by parsing these screenshots using a combination of text analysis and computer vision techniques. We rely on this previous analysis and consider the dataset of 78468 screenshots as our initial dataset. We note that these screenshots exhibit arbitrary structures, as WSB users do not use a specific app and screenshots are handmade, resulting in a lack of standardized format. Through manual inspection of a sample of the screenshots, we identify a set of keywords that proved useful in narrowing down the selection to those representing actual investments (6525 screenshots), as discussed in the main text. Our extraction technique, for instance, does not accommodate the possibility of multiple orders in a single screenshot, as we extract only the highest value.
Supplementary Figure S7: Comparison between OCR and human extraction and stocks rank. a) Scatter plot displaying the comparison between position values extracted using OCR and human visual inspection. b) Treemap rank chart showing the occurrences of stock tickers in screenshots.
Stock A Stock B Number of Screenshots
AMC PLTR 56
GME NIO 62
NOK PLTR 62
NIO PLTR 65
AMC NIO 70
NIO TSLA 80
PLTR TSLA 81
AMC TSLA 98
BB PLTR 132
GME TSLA 148
GME PLTR 183
BB NOK 457
GME NOK 484
AMC BB 648
AMC NOK 698
BB GME 719
AMC GME 1773
@@
Twitter Trading Volume Reddit Date t p-value t p-value t p-value 2020-12-17 -4.296 $0$4.299 $0$4.429 0.000 2020-12-18 -4.251 0.001 -4.236 0.001 -4.504 0.000 2020-12-19 -4.894 $0$4.233 0.001 -4.405 0.000 2020-12-20 -4.894 $0$4.233 0.001 -4.405 0.000 2020-12-21 -4.894 $0$4.233 0.001 -4.405 0.000 $-2$5.000 $0$4.168 0.001 -4.350 0.000 2020-12-23 -4.627 $0$5.939 $0$5.031 0.000 2020-12-24 -5.182 $0$4.987 $0$5.315 0.000 2020-12-25 -5.149 $0$4.134 0.001 -4.947 0.000 2020-12-26 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-27 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-28 -4.853 $0$3.853 0.002 -4.601 0.000 2020-12-29 -4.751 $0$3.880 0.002 -4.666 0.000 2020-12-30 -4.729 $0$4.091 0.001 -4.712 0.000 2020-12-31 -1.841 0.360 -4.280 $0$5.348 0.000 2021-01-01 -3.202 0.020 -4.657 $0$5.688 0.000 2021-01-02 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-03 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-04 -2.931 0.042 -4.244 0.001 -5.259 0.000 2021-01-05 -2.938 0.041 -4.258 0.001 -5.288 0.000 2021-01-06 -2.698 0.074 -4.552 $0$4.940 0.000 2021-01-07 -2.715 0.072 -4.689 $0$4.820 0.000 2021-01-08 -2.788 0.060 -5.738 $0$5.764 0.000 2021-01-09 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-10 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-11 -2.931 0.042 -6.234 $0$6.076 0.000 2021-01-12 -3.698 0.004 -6.108 $0$5.725 0.000 2021-01-13 -3.815 0.003 -5.685 $0$4.988 0.000 2021-01-14 -1.185 0.680 -4.290 $0$4.738 0.000 2021-01-15 -5.310 $0$4.294 $0$4.785 0.000 2021-01-16 -5.680 $0$4.425 $0$5.053 0.000 2021-01-17 -5.680 $0$4.425 $0$5.053 0.000 2021-01-18 -5.680 $0$4.425 $0$5.053 0.000 2021-01-19 -5.346 $0$4.167 0.001 -4.937 0.000 2021-01-20 -5.225 $0$4.147 0.001 -4.752 0.000 2021-01-21 -5.104 $0$3.747 0.004 -4.792 0.000 $-2$5.362 $0$4.006 0.001 -4.843 0.000 2021-01-23 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-24 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-25 -5.393 $0$3.965 0.002 -4.949 0.000 2021-01-26 -4.094 0.001 -4.004 0.001 -5.068 0.000 2021-01-27 -4.972 $0$3.601 0.006 -10.878 0.000 2021-01-28 -4.298 $0$4.008 0.001 -3.058 0.030 2021-01-29 -4.189 0.001 -4.101 0.001 -2.560 0.102 2021-01-30 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-01-31 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-02-01 -3.637 0.005 -4.127 0.001 -2.636 0.086 2021-02-02 -3.254 0.017 -4.352 $0$2.881 0.048 2021-02-03 -3.864 0.002 -4.358 $0$2.767 0.063 2021-02-04 -3.460 0.009 -3.839 0.003 -2.696 0.075 2021-02-05 -3.702 0.004 -3.995 0.001 -2.989 0.036 2021-02-06 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-07 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-08 -3.690 0.004 -4.153 0.001 -3.051 0.030 2021-02-09 -3.001 0.035 -3.754 0.003 -2.262 0.184 2021-02-10 -2.733 0.069 -3.412 0.011 -2.268 0.182 2021-02-11 -2.702 0.074 -3.446 0.009 -2.765 0.063 2021-02-12 -3.432 0.010 -3.196 0.020 -2.123 0.235 2021-02-13 -3.045 0.031 -3.006 0.034 -1.854 0.354